OpenStudy (anonymous):

I am reposting this because I keep ending up with the wrong answer. Can someone my answer? Jerry currently has an account balance of $1,109.80. His initial deposit on the account was $775 and it earned 2.7% simple interest. How long has Jerry held the account? 53 years <-- My answer 37 years 11 years 16 years

OpenStudy (anonymous):

theres an error

OpenStudy (anonymous):

That is exactly how I have it set up.

OpenStudy (anonymous):

where?

OpenStudy (anonymous):

yes. A is actually just the return, and what he has on account - A+P

OpenStudy (anonymous):

so his return is actually just $1,109.80 - P $1,109.80 - $775 =$334.8

OpenStudy (anonymous):

I don't see where you got the return part

OpenStudy (anonymous):

because the formula is actually: I = P r t !!!

OpenStudy (anonymous):

and not A=

OpenStudy (anonymous):

Oh!

OpenStudy (anonymous):

it just calculates the interest

OpenStudy (anonymous):

ok, so the total interest earned is I = A - P

OpenStudy (anonymous):

I = $1,109.80 - $775

OpenStudy (anonymous):

current account balance A - deposit principal P == interest that was earned

OpenStudy (anonymous):

so 334.8 was the interest earned?

OpenStudy (anonymous):

yep :)

OpenStudy (anonymous):

and then we can finally use the simple interest formula I = P r t

OpenStudy (anonymous):

334.8 = P r t

OpenStudy (anonymous):

That makes so much more sense! Can you hold on a sec while I plug in the answer choices again? (:

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

16 years?

OpenStudy (anonymous):

334.8 = P r t 334.8 = $775 r t with r=2.7%/100 = 0.027 334.8 = 775*0.027 * t 334.8 = 20.925 * t |:20.925 16 = t t=16 time - 16 years

OpenStudy (anonymous):

So I was right?! :D

OpenStudy (anonymous):

yes, 16 years should be the answer :)

OpenStudy (anonymous):

Thank you so much! You were very helpful!

OpenStudy (anonymous):

A medal for you! :D

OpenStudy (anonymous):

thx ; )